# Reliable uncertainty with cheaper neural network ensembles: a case study in industrial parts classification

### v1

@@ 5.6 Cost-performance analysis
The plot illustrates that deep and batch ensemble significantly outperform the other models, boasting high accuracy on the ID set and high diversity quality. Batch ensemble, however, is much more computationally efficient than deep ensemble. As such, batch ensemble is a cost-effective and competitive alternative to deep ensemble. Conversely, the snapshot ensemble demonstrates decent diversity quality but has low ID accuracy, failing to match even the single NN. The MIMO ensemble lags behind the other ensembles in both aspects.
@@ 6 Conclusion
Reliable uncertainty estimation in NNs is crucial for their safe deployment in OR, as NNs tend to make confident yet incorrect predictions in OOD scenarios. Deep ensembles have demonstrated robust predictive performance and uncertainty estimates. However, their deployment in practice poses challenges due to substantial computational and memory requirements. This is the first study to provide a comprehensive comparison of a single NN, a deep ensemble, and three efficient NN ensembles. In addition, we propose a Diversity Quality score to quantify the ensembles’ performance on the ID and OOD sets in one single metric. The efficient ensembles include a snapshot ensemble, batch ensemble, and MIMO ensemble, applied on an OR case study in industrial parts classification.
The superior performance of deep and batch ensemble over the single NN highlights the need for reliable uncertainty estimates when deploying NNs in operational settings. Furthermore, batch ensemble emerges as a cost-effective and competitive alternative to deep ensembles. It surpasses the deep ensemble both in terms of accuracy and uncertainty while demonstrating a training time speedup of 7x, a test time speedup of 8x, and 9x memory savings for ensemble size 10. Its robust uncertainty estimates are confirmed by the strong diversity quality score. Additionally, it exhibits similar uncertainty behavior when confronted with OOD data, and selecting the ensemble size is equally straightforward: the more members the better the results.
Several directions for future work are possible. Expanding studies on industrial parts classification to datasets containing more object classes would validate the findings and better simulate real-world industrial scenarios. Furthermore, ensembles hold potential in various other OR domains. For instance, in tasks like household waste sorting or quality control within production plants, the emergence of new product types or failure modes requires reliable uncertainty in classifiers.
@@ Acknowledgments
Reliable uncertainty with cheaper neural network ensembles: a case study in industrial parts classification
Appendix A Results for MIMO with 4 and 10 ensemble members
Table 3 shows the classification accuracy and the NLL metrics on the ID set for both MIMO ensemble configurations, consisting of 4 and 10 ensemble members. The single NN is added for reference. The 10-member ensemble demonstrates superior accuracy compared to its 4-member counterpart, aligning closely with the performance of the single NN.
Table 3: Performance on the ID set (MIMO with 4 and 10 members)
Model Accuracy (%) $\uparrow$ NLL $\downarrow$
Single 96.61 $\pm$ 0.18 0.1121 $\pm$ 0.0039
MIMO E ($M=4$) 95.96 $\pm$ 0.44 0.1168 $\pm$ 0.0092
MIMO E ($M=10$) 96.51 $\pm$ 0.23 0.1258 $\pm$ 0.0088
Figure 11 displays the NRA for both models based on the total uncertainty. The NRA curve of the 10-member ensemble exhibits a notably slower increase compared to the 4-member ensemble. As such, it performs as poorly as the single NN.
Figure 11: Non-rejected accuracy (MIMO with 4 and 10 members)
In conclusion, using the MIMO ensemble with 10 members instead of the optimal size of 3–4 reported by the authors does not seem appropriate. Although the ID accuracy is better than the 4-member configuration, the uncertainty estimates are poor, practically matching the performance of the single NN.
Appendix B Results with 4 ensemble members
This section presents results where all ensembles have 4 members. It is important to highlight that the MIMO ensemble is already configured with 4 members in the main body of the manuscript. This decision was made because using 10 members would restrict the capacity of individual members too much. Therefore, the outcomes presented for the deep, batch, and snapshot ensembles are different from those in the main body.
B.1 In-distribution performance
Table 4 shows the classification accuracy and the NLL metrics on the ID set. The difference in performance between esembles of 4 or 10 members is minimal. The standard error bounds of the accuracy scores intersect across all three ensembles with an adjusted ensemble size. However, batch and snapshot ensembles exhibit marginally higher NLLs. Consequently, the conclusions closely mirror those derived from configurations with 10 ensemble members. In terms of accuracy and NLL, deep ensemble outperforms the single NN, with batch ensemble even surpassing the deep ensemble. In contrast, snapshot and MIMO ensembles lag behind, performing worse than the single NN. Additionally, worth noting that while standard error margins overlap, the snapshot ensemble demonstrates slightly higher accuracy with an ensemble size of 4. Despite this marginal improvement, its performance remains inferior to that of the singular NN.
Table 4: Performance on the ID set (4 ensemble members)
Model Accuracy (%) $\uparrow$ NLL $\downarrow$
Single 96.61 $\pm$ 0.18 0.1121 $\pm$ 0.0039
Deep E ($M=4$) 97.95 $\pm$ 0.09 0.0748 $\pm$ 0.0024
Snapshot E ($M=4$) 94.97 $\pm$ 0.25 0.1625 $\pm$ 0.0051
Batch E ($M=4$) 98.60 $\pm$ 0.17 0.0586 $\pm$ 0.0078
MIMO E ($M=4$) 95.96 $\pm$ 0.44 0.1168 $\pm$ 0.0092
B.2 Uncertainty density
Figure 12 depicts the density distributions of total, data, and model uncertainty. Notably, the uncertainty levels on the ID set remain consistent with the configuration employing 10 members. Conversely, the uncertainty on the OOD set slightly lower across all three ensembles following adjustments to the ensemble size.
(a) Total uncertainty
(b) Data uncertainty
(c) Model uncertainty
Figure 12: Densities for the data, model, and total uncertainty on the ID and OOD set (4 ensemble members)
B.3 Classification with rejection
@@
Figure 13 displays the NRA for all models based on the total uncertainty. We observe a sizable decrease in NRA when transitioning from 10 ensemble members to only 4. Specifically, at lower uncertainty thresholds, both snapshot and deep ensembles exhibit a decrease of 10 percentage points, while the batch ensemble experiences a more pronounced decrease of 15 percentage points. Batch ensemble seems to be affected more by the smaller ensemble size than the other two ensembles. Despite reaching the same NRA than MIMO at low uncertainty thresholds, the batch ensemble demonstrates consistently superior performance for moderate to larger thresholds.
Figure 13: Non-rejected accuracy (4 ensemble members)
B.4 Diversity analysis
Figure 13(b) displays a scatterplot of diversity scores on the ID and OOD test set, where each point represents an ensemble member. The observations closely resemble those of the configuration with 10 members. However, on the OOD set, the deep ensemble is now slightly more diverse than the batch ensemble. Overall, the deep and batch ensemble are still superior over the other ensembles.
(a) ID set
(b) OOD set
Figure 14: Diversity on the ID and OOD set (4 ensemble members). Each point represents the diversity of a trained model against the model averaged over all members
Figure 15 shows the $DQ_{1}$-scores, where each point represents an ensemble member. The observations are in line with the configuration with 10 members and follow closely on the diversity analysis in section B.4. The deep ensemble members have slightly higher $DQ_{1}$-scores than the batch ensemble, both still outperforming the snapshot and MIMO ensemble.
Figure 15: $DQ_{1}$-scores (4 ensemble members). Each point represents the $DQ_{1}$-score of an ensemble member
B.5 Computational cost
Figure 16 depicts the computational costs of the ensembling techniques. The values are normalized relative to the single model. As anticipated, the computational cost of the deep ensemble maintains a linear relationship with the ensemble size. Additionally, the training time of the snapshot ensemble remains consistent with that of the single NN, while its evaluation time and memory cost align with those of the deep ensemble. Conversely, the costs associated with the batch and MIMO ensembles show minimal reduction, resulting in costs that are only marginally higher than those of the single NN. Consequently, using a batch ensemble with an increased ensemble size is an interesting option, as performance typically improves with larger ensemble sizes.
(a) Training cost
(b) Evaluation cost
(c) Memory cost
Figure 16: Training, evaluation, and memory cost (4 ensemble members)
B.6 Cost-performance analysis
@@
Figure 17 displays a bubble chart combining the predictive accuracy, diversity scores, and computational costs of all models described in previous sections. The observations closely resemble those of the configuration with 10 members. Deep and batch ensemble significantly outperform the other models, showcasing much higher ID accuracy and OOD diversity. Batch ensemble is still substantially more efficient than deep ensemble and is an excellent alternative.

### v2

@@ 6.6 Cost-performance analysis
The plot illustrates that deep and batch ensemble significantly outperform the other models, boasting high accuracy on the ID set and high Diversity Quality. Batch ensemble, however, is much more computationally efficient than deep ensemble. It achieves strong performance with few resources by sharing information between the ensemble members, unlike a deep ensemble consisting of entirely independent members. Furthermore, it is relatively easy to optimize as more ensemble members generally translate to better performance, unlike snapshot and MIMO ensembles. As such, batch ensemble is a cost-effective and competitive alternative to deep ensemble. Conversely, the MIMO ensemble demonstrates decent Diversity Quality but has low ID accuracy, failing to match even the single NN. The snapshot ensemble lags behind the other ensembles in both aspects.
@@ 7 Conclusion
NN image classifiers deployed in industrial settings are likely to be confronted with OOD scenarios, where they often make confident yet incorrect predictions. Deep ensembles have been shown to overcome these challenges by quantifying uncertainty in their own predictions but are computationally expensive. This study therefore investigates more efficient NN ensembles in order to obtain uncertainty estimates at a lower cost, applied on a dataset of industrial parts classification. It is the first to provide a comprehensive comparison among the methods and it proposes a novel Diversity Quality score to quantify the ensembles’ performance on the ID and OOD sets in one single metric.
The superior performance of deep and batch ensemble over the single NN highlights the need for reliable uncertainty estimation to safely deploy NNs in operational settings. By sharing information between the ensemble members, batch ensemble emerges as a cost-effective and competitive alternative to deep ensembles. It matches the deep ensemble both in terms of accuracy and Diversity Quality while demonstrating savings in training time, test time, and memory storage. Additionally, it exhibits similar uncertainty behavior when confronted with OOD data, and selecting the ensemble size is equally straightforward: more members generally results in better performance.
Several directions for future work are possible. Expanding studies on industrial parts classification to datasets containing more object classes would validate the findings and better simulate real-world industrial scenarios. Furthermore, ensembles hold potential in various other OR domains. For instance, in tasks like household waste sorting or quality control within production plants, the emergence of new product types or failure modes requires reliable uncertainty in classifiers. Finally, it is promising to investigate sim-to-real setups (Tobin et al., 2017) where the NN is only trained on synthetic images and evaluated on real images, requiring less collecting and labeling of real images. As such, this workflow exhibits an inherent distribution shift, also requiring reliable uncertainty quantification.
@@ Acknowledgments
Fast and reliable uncertainty quantification with neural network ensembles for industrial image classification
Appendix A Dataset
Table 5 displays a detailed overview of the dataset, per use case and object.
Table 5: Number of images per use case and per object
Set Use case Object Train Validation Test
ID Use case 1 Airgun 1020 180 300
Electricity12 1020 180 300
Hammer 1020 180 300
Hook 1020 180 300
Plug 1020 180 300
OOD Use case 2 Fork1 1020 180 300
Fork2 1020 180 300
Fork3 1020 180 300
Use case 3 CouplingHalf 300
Gear1 300
Gear2 300
Pinion 300
Use case 4 Cross 300
Pin1 300
Pin2 300
Appendix B Hyperparameter Tuning
All models are trained using the Adam optimizer and the snapshot ensemble uses the custom cyclic learning rate schedule. The single NN, deep ensemble, and snapshot ensemble are trained for 200 epochs. Following Nado et al. (2021), the batch ensemble and MIMO ensemble are trained for 50% longer (i.e., 300 epochs) because they take longer to converge. The training batch size was set to 512. For batch ensemble, we initialize the fast weights to be random sign vectors, as the authors of batch ensemble note that this encourages diversity among the members.
Hyperparameter tuning is used to select the initial learning rate and the L2 regularization weight. For the batch ensemble and MIMO ensemble, we tune one additional hyperparameter. Hyperparameters are tuned independently for each ensemble type, except for the deep ensemble as it simply consists of multiple tuned single NNs. Table 6 shows the hyperparameter settings. Hyperparameters are optimized through a Tree-structured Parzen Estimator using Optuna (Akiba et al., 2019) for 20 trials. We follow Ovadia et al. (2019) and minimize the NLL on the validation set; they optimize NLL rather than accuracy since the former is a proper scoring rule. Table 7 shows the optimal validation NLL obtained during hyperparameter tuning. In the main results section, we work with a snapshot ensemble of $M=8$ members and a MIMO ensemble of $M=3$ members because this yields the lowest validation NLL for their ensemble type.
Table 6: Hyperparameter values
Hyperparameter Range Selected value
Single
Epochs {200} 200
[…]
Snapshot ensemble ($M=4$)
Epochs {200} 200
[…]
Snapshot ensemble ($M=6$)
Epochs {200} 200
[…]
Snapshot ensemble ($M=8$)
Epochs {200} 200
[…]
Batch ensemble ($M=4$)
Epochs {300} 300
[…]
Fast weight initialisation {“random sign”} “random sign”
Fast weight LR multiplier [0.1, 1] $0.490,045,542,908,698,3$
Batch ensemble ($M=8$)
Epochs {300} 300
[…]
Fast weight initialisation {“random sign”} “random sign”
Fast weight LR multiplier [0.1, 1] $0.389,610,642,499,936,2$
MIMO ensemble ($M=3$)
Epochs {300} 300
[…]
Input repetition probability [0.0, 0.8] 0.3
MIMO ensemble ($M=4$)
Epochs {300} 300
[…]
Input repetition probability [0.0, 0.8] 0.8
Table 7: Optimal validation NLL
Model Ensemble size $M$ Val NLL
Single 0.1040
Snapshot ensemble 4 0.1119
6 0.1178
8 0.1042
Batch ensemble 4 0.0460
8 0.0494
MIMO ensemble 3 0.1023
4 0.1255
Appendix C Results for all ensemble configurations
This section presents results for all ensemble configurations. The ensemble results reported in the main body are replicated to provide easier comparisons.
Table 8 shows the classification accuracy and the NLL metrics on the ID set. For deep and batch ensemble, the difference between ensemble sizes 4 or 8 is minimal. For snapshot and MIMO ensemble, the configuration selected based on the validation NLL ($M=8$ for snapshot and $M=3$ for MIMO) is on par or better than the other ensemble sizes.
Table 8: Performance on the ID set (all configurations)
Model Ensemble size $M$ Accuracy (%) $\uparrow$ NLL $\downarrow$
Single 97.72 $\pm$ 0.14 0.1047 $\pm$ 0.0089
Deep ensemble 4 98.51 $\pm$ 0.09 0.0496 $\pm$ 0.0027
8 98.53 $\pm$ 0.09 0.0440 $\pm$ 0.0011
Snapshot ensemble 4 96.68 $\pm$ 0.26 0.1029 $\pm$ 0.0065
6 95.20 $\pm$ 0.16 0.1526 $\pm$ 0.0030
8 97.15 $\pm$ 0.10 0.1048 $\pm$ 0.0031
Batch ensemble 4 98.47 $\pm$ 0.07 0.0514 $\pm$ 0.0026
8 98.65 $\pm$ 0.05 0.0504 $\pm$ 0.0035
MIMO ensemble 3 96.63 $\pm$ 0.30 0.1057 $\pm$ 0.0061
4 96.57 $\pm$ 0.16 0.1085 $\pm$ 0.0020
@@ Appendix C Results for all ensemble configurations
Figure 12 displays the NRA for all ensemble configurations based on the total uncertainty. For deep and batch ensemble, we observe a sizable decrease in NRA when transitioning from 8 ensemble members to only 4. Specifically, at lower uncertainty thresholds, the NRA curves with $M=4$ are 5–8 percentage points lower than those with $M=8$. This finding shows that although configurations with $M=4$ show similar ID performance than with $M=8$, the larger number of members yields better performance on the OOD set. For snapshot ensemble, the selected configuration $M=8$ performs on par or better than the other ensemble sizes. The selected MIMO configuration $M=3$ outperforms the other ensemble size $M=4$ by a significant margin.
Figure 12: Non-rejected accuracy (all configurations)
Figure 13 displays a bubble chart combining the predictive accuracy, diversity scores, and computational costs of ensemble configurations. Similarly to the $M=8$ configuration, for $M=4$ batch ensemble exhibits comparable performance to the deep ensemble, despite being much less computationally expensive. For snapshot ensemble, the alternative configurations with $M=4$ and $M=6$ are inferior to the selected configuration $M=8$ as the $DQ_{1}$-scores are equal but the accuracy values are lower. The alternative MIMO configuration with $M=4$ is inferior to the selected $M=3$ configuration as the accuracy is equal but the $DQ_{1}$-score is much worse.
Figure 13: Accuracy on the ID set and $DQ_{1}$-score (all configurations). Each point represents an ensemble model; the point size represents the weighted computational cost relative to a single NN. Models in the upper-right corner perform best
@@
As such, we conclude that deep and batch ensemble significantly outperform the other models, showcasing much higher ID accuracy and $DQ_{1}$-scores. Batch ensemble is substantially more efficient than deep ensemble and forms an excellent alternative.
Appendix D $DQ_{\beta}$-score sensitivity
Figure 14 displays a bubble chart of the $DQ_{\beta}$-scores for varying $\beta$ values, where the bubble size denotes the computational expense as discussed in Subsection 6.6. Subfigure 14(a) shows the $DQ_{1}$-score, which corresponds to the vertical axis of Figure 11.
Subfigure 14(b) shows the results for $\beta=0.25$, indicating that the ID performance is 4 times more important to the practitioner than the OOD performance. The bubbles are now closer to each other because the largest performance differences are observed on the OOD set, which is now weighted less in the $DQ$-score. Note that the $DQ$-scores are now substantially higher as the IDD scores are higher than the OODD scores.
Subfigure 14(c) shows the results for the parameter $\beta=4$, reflecting the practitioner’s priority for OOD performance over ID performance. In contrast to before, the bubbles now move further apart because we observe larger differences in OODD. For example, the $DQ$-score of the snapshot ensemble is now lower relative to the deep ensemble as the snapshot ensemble performed poorly in terms of OODD.
The ranking of the methods remains unchanged, but the relative differences between them do shift. The $DQ$-score helps practitioners more easily assess whether the potential improvement in diversity performance justifies the development of a more computationally intensive model. Management can set the $\beta$ value, while the machine learning team can then choose the most suitable model type for production.
(a) $DQ_{1}$-score
(b) $DQ_{0.25}$-score
(c) $DQ_{4}$-score
Figure 14: $DQ_{\beta}$-scores for varying $\beta$ values